OpenStudy (love10129151):

4. What is the solution of the system? y = −3x 3x + 2y = 6(1 point) (2, 6) (−1, 3) (−2, 6) (1.5, −4.5)

OpenStudy (jdoe0001):

y = −3x <----- notice what "y" equals to, so use that and substitute in the 2nd one 3x + 2y = 6 what would that look like?

OpenStudy (love10129151):

you look at three

OpenStudy (jdoe0001):

at three?

OpenStudy (love10129151):

no

OpenStudy (love10129151):

or yes

OpenStudy (jdoe0001):

what do you mean?

OpenStudy (love10129151):

you look at three or no

OpenStudy (jdoe0001):

\(\bf y = −3x\\ 3x + 2y = 6\qquad \textit{substituting "y" in the 2nd one}\\ \quad \\ \color{red}{y} = -3x\quad thus \quad 3x + 2y = 6\implies 3x + 2(\color{red}{-3x}) = 6\) so solving for "x", what would you get?

OpenStudy (love10129151):

ooo ok

OpenStudy (love10129151):

so the answer is ...........

OpenStudy (jdoe0001):

... what did you get for "x" ?

OpenStudy (love10129151):

so x is 2

OpenStudy (love10129151):

i am i right or no

OpenStudy (jdoe0001):

well \(\bf 3x+2(-3x)=6\implies 3x-6x=6\implies -3x=6\implies x=\cfrac{6}{-3}\\ \quad \\ x=-2\) thus what's "y"? well, we know that y = -3x, so since x = -2, then y = -3(-2)

OpenStudy (love10129151):

so its (-2,6)

OpenStudy (jdoe0001):

yeap

OpenStudy (love10129151):

thank can you help me with one more

OpenStudy (jdoe0001):

you can always post anew